# A Cold and Super-Puffy Planet on a Orbit

### v2

@@ Abstract
We report the discovery of TOI-4507 b, a transiting sub-Saturn with a density on a 105-day orbit around a 700 Myr old F star. The transits were detected using data from TESS as well as the Antarctic telescope ASTEP. A joint analysis of the light curves and radial velocities from HARPS, FEROS, and CORALIE confirmed the planetary nature of the signal by limiting the mass to be below at 95% confidence. The radial velocities also exhibit the Rossiter-McLaughlin effect and imply that the With these characteristics, TOI-4507 is one of the longest-period systems for which the stellar obliquity has been measured, and the planet is among the longest-period and youngest “super-puff” planets yet discovered.
@@
1 Introduction
6,000 exoplanets have been discovered [^1] https://exoplanetarchive.ipac.caltech.edu/, most of them using the methods of transits and radial velocities (RVs). The Kepler (W. J. Borucki et al., 2010) and Transiting Exoplanet Survey Satellite (TESS; G. R. Ricker et al., 2015) discoveries have revealed valuable information about the demographics of short-period exoplanets (see, e.g., a review by W. Zhu & S. Dong, 2021). However, because the probability of observing transits decreases with orbital distance, it is challenging to explore the outer regions of planetary systems using transits (e.g., B. S. Gaudi et al., 2005). Likewise, RV surveys have revealed valuable information about the demographics of cold Jupiters (e.g., A. Cumming et al., 2008; R. A. Wittenmyer et al., 2016; R. B. Fernandes et al., 2019; A. S. Bonomo et al., 2023), but because the RV semiamplitude decreases with decreasing planetary mass and increasing semimajor axis, it has been difficult to study less massive cold worlds.
@@ 1 Introduction
In this letter, we present the discovery and characterization of TOI-4507 b, a cold planet transiting a young F star with an orbital period of 105 days. Our observations reveal that the planet has an unusually low density, making it one of the longest-period super-puffs known to date. Furthermore, our analysis of the Rossiter-McLaughlin (RM) effect (R. Rossiter, 1924; D. McLaughlin, 1924) shows that the planet has a A more detailed validation of TOI-4507 b using TESS and ASTEP data appears in a separate paper (E. Rea et al., 2025).
2 Observations
2.1 Photometry
2.1.1 TESS
@@ 2.2.1 HARPS
The spectra were reduced Because of the warm-up of HARPS on March 23, 2020, due to the COVID-19 pandemic, the data taken before and after this date were considered as having come from different instruments (HARPS1 and HARPS2) with different additive RV offsets and “jitter” parameters (the level of uncertainty in excess of the formal uncertainty). The HARPS RVs of the target, together with the best model, are shown in Figure 1
2.2.2 FEROS
We also observed the target using the Fiber-fed Extended Range Optical Spectrograph (FEROS; A. Kaufer et al., 1999) installed in the MPG/ESO 2.2 m telescope in La Silla Observatory in Chile. FEROS is an echelle spectrograph with a resolving power of $R\approx 48,\!000$ and a wavelength range coverage of 350920350-920 nm. We obtained a total of 11 out-of-transit spectra of the star between August 2019 and November 2023. The exposure times range between 600 s and 1800 s. FEROS observations were also processed using the ceres pipeline and RVs extracted using the CCF method. The median S/N of the FEROS spectra is 95 per resolution element at 550 nm, with a median RV uncertainty of 9.5 m/s. The resulting FEROS RVs for TOI-4507, along with the best-fit model, are presented in Figure 1
2.2.3 CORALIE
We also observed the target using the CORALIE spectrograph (D. Queloz et al., 2001) installed in the Swiss 1.2 m Leonhard Euler Telescope at ESO’s La Silla Observatory. CORALIE is an echelle spectrograph with a resolving power of $R\approx 60,\!000$, which covers the wavelength range of 390680390-680 nm. We observed TOI-4507 using exposure times varying between 900 and 1800 s, together with the simultaneous Fabry-Pérot étalon. We obtained a total of 10 out-of-transit spectra between October 2019 and February 2021. Data reduction was done using the dedicated CORALIE data reduction pipeline, which also derives RV using the CCF technique and a G2 mask. CORALIE observations have a median RV uncertainty of 28 m/s and are shown in Figure 1 along with the best model.
2.3 High-Resolution Imaging
@@ 4 Photometric analysis
To determine the transit ephemerides of TOI-4507 b and look for transit timing variations (TTVs), we analyzed the photometry presented in Section 2 with the juliet code (N. Espinoza et al., 2019). This code uses batman (L. Kreidberg, 2015) for the transit model and the dynesty dynamic nested sampler (J. S. Speagle, 2020) to sample the posteriors. We placed uniform priors on the impact parameter $b$ and radius ratio $R_{p}/R_{\star}$, with a Gaussian prior on the stellar density $\rho_{\star}$ that was constrained in Section 3. We sampled the limb darkening parameters using the quadratic $q_{1}$ and $q_{2}$ parameters from D. M. Kipping (2013) with uniform priors. We placed Gaussian priors for each transit mid-point based on the expected values calculated from the orbital period and time of mid-transit from the SPOC solution, placing a large width of 1 day on the prior to not impact the derived transit midpoints. To account for variability and systematic noise in the TESS light curve, we included a Matern-3/2 Gaussian Process (GP) as implemented in celerite (D. Foreman-Mackey et al., 2017) and available in juliet. From this analysis, we found $P=104.61573\pm 0.00008$ d and $t_{0}=2458413.8572\pm 0.0007$ (BJD). The nearly continuous TESS time series between Sectors 5 and 13, 32 and 39, and 61 and 69 left no doubt that the true orbital period is about 105 days, as consecutive transits are seen in the TESS data. Here, we also obtained a detrended TESS light curve that was used in the Finally, we also found tentative evidence for TTVs, as the transit observed with ASTEP in 2025 deviates by $\sim 25$ minutes from the linear ephemerides solution. The rest of the transits do not show this behavior, being in all cases within $\sim 5$ minutes of the linear solution.
5 Joint Photometric and RV Modeling
6 Rossiter-McLaughlin Effect
@@
$t_{0}$ Transit midpoint (BJD - 2,458,000) $\mathcal{U}(413.8472,413.8612)^{a}$ $2458413.8515\pm 0.0008$
@@
$\gamma_{\rm HARPS2}$ HARPS2 RV offset (m/s) $\mathcal{U}(-100,100)$ $26095_{-10}^{+9}$
@@
7 Discussion
7.1 TOI-4507 b in context
TOI-4507 b appears to be an interesting planet in several ways. Figure 4 shows the results for the projected obliquity $\lambda$ along with those of other systems. The obliquity data were drawn from TEPCat [^2] https://www.astro.keele.ac.uk/jkt/tepcat/ (J. Southworth, 2011) and the age data were drawn from the NASA Exoplanet Archive (R. L. Akeson et al., 2013; J. L. Christiansen et al., 2025), both as of May 2025. Evidently, after HIP 41378 d (S. Grouffal et al., 2022), TIC 241249530 b (A. F. Gupta et al., 2024), and HD 80606 b (e.g., G. Hébrard et al., 2010). Here, we define Neptunes/sub-Saturns as the class of planets with $10<M_{p}/M_{\oplus}<90$, with the subclass of super-puffs having $\rho_{p}<0.3$ g/cm3.
TOI-4507 b joins a growing population of Neptune/sub-Saturn
@@ 7.1 TOI-4507 b in context
Additionally, with an age of $\sim 700$ Myr, TOI-4507 b is the youngest super-puff with an obliquity measurement. Further, it is one of the youngest Neptunes/sub-Saturns
@@ 7.2 Search for companions
If the RV trend is produced by a planetary companion, it should have an orbital period $P>5.6$ yr, which corresponds to the baseline of the current RV observations. The parameters of this possible companion could be constrained with additional RV and/or Gaia DR4 and DR5 astrometric measurements. When removing the signal of TOI-4507 b and the RV trend from the data, we do not see clear additional signals in the RV data, which rules out the presence of planets producing similar or larger RV signals than TOI-4507 b.
@@
7.3 Explaining the Inferred Low Density
@@
8 Conclusions
We have reported the discovery and characterization of TOI-4507 b, a transiting cold Neptune on a orbit. With a radius of $8.2\pm 0.1\,R_{\oplus}$, a mass $20\,M_{\oplus}$, and a bulk density <0.2<0.2 g/cm3, TOI-4507 b is among the longest-period ($P\sim 105$ d) super-puffs discovered to date.
@@ 8 Conclusions
TOI-4507 thus joins a small but growing population of providing a valuable laboratory for understanding the formation, evolution, and diversity of planetary systems.
@@
We would like to thank Yubo Su for suggesting the rotational constraint on the sizes of the rings. We also thank Songhu Wang, Xian-Yu Wang, and the Red Worlds Lab group at the University of Amsterdam for their useful discussions and suggestions that strengthened this manuscript. J.I.E.-R. gratefully acknowledges support from the John and A-Lan Reynolds Faculty Research Fund, from the ANID BASAL project FB210003, and from the ANID Doctorado Nacional grant 2021-21212378. R.B. acknowledges support from FONDECYT Project 1241963 and from ANID – Millennium Science Initiative – ICN12_009. Support for M.C. is provided by ANID’s Millennium Science Initiative through grants ICN12_12009 and AIM23-001, awarded to the Millennium Institute of Astrophysics (MAS); by ANID/FONDECYT Regular grant 1231637; and by ANID/Basal (CATA) grant FB21003. $C.$ acknowledges support from ANID through FONDECYT post-doctoral grant number 3230283. A.V.F. acknowledges the support of the Institute of Physics through the Bell Burnell Graduate Scholarship Fund. This paper was based on observations collected at the European Southern Observatory under ESO programmes 0104.C-0413(A), 106.21ER.001, 112.25W1.001, 114.27CS.001, and 115.286G.001. This paper also made use of data collected by the TESS mission and are publicly available from the Mikulski Archive for Space Telescopes (MAST) operated by the Space Telescope Science Institute (STScI). Funding for the TESS mission is provided by NASA’s Science Mission Directorate. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center for the production of the SPOC data products. This work makes use of observations from the ASTEP telescope. ASTEP benefited from the support of the French and Italian polar agencies IPEV and PNRA in the framework of the Concordia station program and from OCA, INSU, Idex UCAJEDI (ANR- 15-IDEX-01) and ESA through the Science Faculty of the European Space Research and Technology Centre (ESTEC). This research also received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 803193/BEBOP), from the Science and Technology Facilities Council (STFC; grant No. ST/S00193X/1, ST/W002582/1 and ST/Y001710/1), and from the ERC/UKRI Frontier Research Guarantee programme (CandY/ EP/Z000327/1). This work has been carried out within the framework of the National Centre of Competence in Research PlanetS supported by the Swiss National Science Foundation under grants 51NF40_182901 and 51NF40_205606. This paper is also based on observations obtained at the SOAR telescope, which is a joint project of the Ministério da Ciência, Tecnologia e Inovações (MCTI/LNA) do Brasil, the US National Science Foundation’s NOIRLab, the University of North Carolina at Chapel Hill (UNC), and Michigan State University (MSU). This research has made use of the Exoplanet Follow-up Observation Program (ExoFOP) and NASA Exoplanet Archive websites, which are operated by the California Institute of Technology under contract with NASA under the Exoplanet Exploration Program.

### v3

# A Cold and Super-Puffy Planet on a Prograde Orbit

@@ Abstract
We report the discovery of TOI-4507 b, a transiting sub-Saturn with a density << 0.2 g/cm3 on a 105-day prograde orbit around a 700 Myr old F star. The transits were detected using data from TESS as well as the Antarctic telescope ASTEP. A joint analysis of the light curves and radial velocities from HARPS, FEROS, and CORALIE confirmed the planetary nature of the signal by limiting the mass to be below $20\,M_{\oplus}$ at 95% confidence. The radial velocities also exhibit the Rossiter-McLaughlin effect and imply that the planet orbits the star in a prograde orbit with a sky-projected obliquity $\lambda=-15_{-44}^{+50}$ deg ($|\lambda|<80$ deg at $3\sigma$). With these characteristics, TOI-4507 is one of the longest-period systems for which the stellar obliquity has been measured, and the planet is among the longest-period and youngest “super-puff” planets yet discovered.
@@
I Introduction
More than 6,000 exoplanets have been discovered [^1] https://exoplanetarchive.ipac.caltech.edu/, most of them using the methods of transits and radial velocities (RVs). The Kepler (W. J. Borucki et al., 2010) and Transiting Exoplanet Survey Satellite (TESS; G. R. Ricker et al., 2015) discoveries have revealed valuable information about the demographics of short-period exoplanets (see, e.g., a review by W. Zhu & S. Dong, 2021). However, because the probability of observing transits decreases with orbital distance, it is challenging to explore the outer regions of planetary systems using transits (e.g., B. S. Gaudi et al., 2005). Likewise, RV surveys have revealed valuable information about the demographics of cold Jupiters (e.g., A. Cumming et al., 2008; R. A. Wittenmyer et al., 2016; R. B. Fernandes et al., 2019; A. S. Bonomo et al., 2023), but because the RV semiamplitude decreases with decreasing planetary mass and increasing semimajor axis, it has been difficult to study less massive cold worlds.
@@ I Introduction
In this letter, we present the discovery and characterization of TOI-4507 b, a cold planet transiting a young F star with an orbital period of 105 days. Our observations reveal that the planet has an unusually low density, making it one of the longest-period super-puffs known to date. Furthermore, our analysis of the Rossiter-McLaughlin (RM) effect (R. Rossiter, 1924; D. McLaughlin, 1924) shows that the planet has a prograde orbit, making it one of the longest-period systems for which the stellar obliquity has been measured. A more detailed validation of TOI-4507 b using TESS and ASTEP data appears in a separate paper (E. Rea et al., 2025).
II Observations
II.1 Photometry
II.1.1 TESS
@@ II.2.1 HARPS
The spectra were reduced using the ceres pipeline (R. Brahm et al., 2017a), which, starting from the raw images, performs all reduction steps and derives precise RVs via the cross-correlation function (CCF) technique. The median signal-to-noise ratio (S/N) of the processed spectra is 50 per resolution element at 550 nm, leading to a median formal RV uncertainty of 4.7 m/s. Because of the warm-up of HARPS on March 23, 2020, due to the COVID-19 pandemic, the data taken before and after this date were considered as having come from different instruments (HARPS1 and HARPS2) with different additive RV offsets and “jitter” parameters (the level of uncertainty in excess of the formal uncertainty). The HARPS RVs of the target, together with the best model, are shown in Figure 1 and available in Appendix A.
II.2.2 FEROS
We also observed the target using the Fiber-fed Extended Range Optical Spectrograph (FEROS; A. Kaufer et al., 1999) installed in the MPG/ESO 2.2 m telescope in La Silla Observatory in Chile. FEROS is an echelle spectrograph with a resolving power of $R\approx 48,\!000$ and a wavelength range coverage of 350920350-920 nm. We obtained a total of 11 out-of-transit spectra of the star between August 2019 and November 2023. The exposure times range between 600 s and 1800 s. FEROS observations were also processed using the ceres pipeline and RVs extracted using the CCF method. The median S/N of the FEROS spectra is 95 per resolution element at 550 nm, with a median RV uncertainty of 9.5 m/s. The resulting FEROS RVs for TOI-4507, along with the best-fit model, are presented in Figure 1 and available in Appendix A.
II.2.3 CORALIE
We also observed the target using the CORALIE spectrograph (D. Queloz et al., 2001) installed in the Swiss 1.2 m Leonhard Euler Telescope at ESO’s La Silla Observatory. CORALIE is an echelle spectrograph with a resolving power of $R\approx 60,\!000$, which covers the wavelength range of 390680390-680 nm. We observed TOI-4507 using exposure times varying between 900 and 1800 s, together with the simultaneous Fabry-Pérot étalon. We obtained a total of 10 out-of-transit spectra between October 2019 and February 2021. Data reduction was done using the dedicated CORALIE data reduction pipeline, which also derives RV using the CCF technique and a G2 mask. CORALIE observations have a median RV uncertainty of 28 m/s and are shown in Figure 1 along with the best model. Data is available in Appendix A.
II.3 High-Resolution Imaging
@@ IV Photometric analysis
To determine the transit ephemerides of TOI-4507 b and look for transit timing variations (TTVs), we analyzed the photometry presented in Section II with the juliet code (N. Espinoza et al., 2019). This code uses batman (L. Kreidberg, 2015) for the transit model and the dynesty dynamic nested sampler (J. S. Speagle, 2020) to sample the posteriors. We placed uniform priors on the impact parameter $b$ and radius ratio $R_{p}/R_{\star}$, with a Gaussian prior on the stellar density $\rho_{\star}$ that was constrained in Section III. We sampled the limb darkening parameters using the quadratic $q_{1}$ and $q_{2}$ parameters from D. M. Kipping (2013) with uniform priors. We placed Gaussian priors for each transit mid-point based on the expected values calculated from the orbital period and time of mid-transit from the SPOC solution, placing a large width of 1 day on the prior to not impact the derived transit midpoints. To account for variability and systematic noise in the TESS light curve, we included a Matern-3/2 Gaussian Process (GP) as implemented in celerite (D. Foreman-Mackey et al., 2017) and available in juliet. From this analysis, we found $P=104.61573\pm 0.00008$ d and $t_{0}=2458413.8572\pm 0.0007$ (BJD). The nearly continuous TESS time series between Sectors 5 and 13, 32 and 39, and 61 and 69 left no doubt that the true orbital period is about 105 days, as consecutive transits are seen in the TESS data. Here, we also obtained a detrended TESS light curve that was used in the subsequent analyses. Finally, we also found tentative evidence for TTVs, as the transit observed with ASTEP in 2025 deviates by $\sim 25$ minutes from the linear ephemerides solution. The rest of the transits do not show this behavior, being in all cases within $\sim 5$ minutes of the linear solution.
The TESS light curve exhibits a quasi-periodic modulation with a periodicity of $\sim 1.7$ days. If this were due to stellar rotation, then the combination of the measurements of rotation period, stellar radius, and projected rotation velocity would imply a low value of $\sin i_{\star}\approx 0$ and a nearly-polar orbit for the planet. This was our initial interpretation. However, a more detailed analysis revealed that the 1.7-day variability comes from a blended variable star that lies in the background (see Appendix B). Without an estimate of the rotation period of TOI-4507, we cannot constrain the inclination of the stellar rotation axis.
V Joint Photometric and RV Modeling
In order to derive the parameters of the planet, we performed a joint analysis of all the available photometry and out-of-transit RVs using the ironman [^2] https://github.com/jiespinozar/ironman code (J. I. Espinoza-Retamal et al., 2023b, 2024). In brief, ironman is a python code that can jointly fit in-transit and out-of-transit RVs together with transit photometry. To model the photometry and RVs, ironman uses radvel (B. J. Fulton et al., 2018) and batman (L. Kreidberg, 2015), respectively. Finally, to get the posteriors ironman uses the dynesty dynamic nested sampler (J. S. Speagle, 2020). A copy of ironman version 1.1.0, which was used in the analysis, is available in Zenodo (see J. I. Espinoza-Retamal, 2025).
To reduce the computational cost, we only considered the detrended TESS data within 7.5 hours of a transit midpoint. We included independent jitter terms for each RV and photometric instrument to account for possible instrumental systematics. We placed uniform priors on all of the parameters except for the stellar density, for which we used the constraint described in Section III. Finally, to account for the possible non-detection of the Keplerian signal in the RVs, we allowed the semiamplitude to take negative values. All priors and resulting posteriors are shown in Table 2.
[…]
There is a hint of a small eccentricity, $e=0.09^{+0.26}_{-0.06}$. Indeed, we fitted the data using both a circular and an eccentric model. For the eccentric case, we sampled the eccentricity and argument of periastron as $\sqrt{e}\cos{\omega}$ and $\sqrt{e}\sin{\omega}$. The Bayesian evidence difference ($\Delta\log{Z}$), being <2<2, is not enough to favor any of the models in particular. As both models returned a fully consistent set of parameters, we elected to formally adopt the values from the eccentric fit to highlight the possible range of eccentricities compatible with the data. This possible eccentricity can also be further constrained with additional RV measurements.
We found evidence for a long-term RV trend, possibly caused by a longer-period outer companion. We modeled the trend using a quadratic model, $\ddot{\gamma}(t-t_{a})^{2}+\dot{\gamma}(t-t_{a})$, where $t_{a}$ was arbitrarily chosen to be the time of the earliest precise RV measurement. The fit gave $\ddot{\gamma}=0.0180\pm 0.0035$ m/s/day/yr and $\dot{\gamma}=-0.12\pm 0.02$ m/s/day, implying a $5\sigma$ detection. Further information about this possible companion, including its orbital inclination, might be possible by combining additional RVs and Gaia DR4 astrometry (e.g., J. I. Espinoza-Retamal et al., 2023a), which would help to understand the system’s dynamical history.
VI Rossiter-McLaughlin Effect
@@
In order to constrain the stellar obliquity, we analyzed the HARPS observations of the RM effect using ironman. To model the RM effect, ironman uses rmfit (G. Stefansson et al., 2020, 2022), which implements the equations from T. Hirano et al. (2010). Given the lack of an estimation of the rotational period of TOI-4507 (see Appendix B), here we only derived the sky-projected obliquity by directly sampling $\lambda$ and $v\sin{i_{\star}}$, both with uninformative uniform priors. We placed informative priors for almost all the rest of the parameters based on the values reported in Table 2. We also placed an informative prior on $\beta$, the intrinsic linewidth accounting for instrumental and macroturbulence broadening. We considered an instrumental broadening of 2.6 km/s because of the HARPS resolution, and for the macroturbulence broadening, we followed S. Albrecht et al. (2012) and used the macroturbulence law for hot stars from D. F. Gray (1984) considering $T_{\rm eff}=6260$ K, which resulted in 5.5 km/s. We added the instrumental and macroturbulence broadening in quadrature to set our prior, with an uncertainty of 2 km/s. Finally, to account for possible RV trends in the RM data due to orbital motion or stellar activity, we allowed for an arbitrary RV slope in the model. All priors and resulting posteriors are shown in Table 3.
As shown in Figure 3, we found that TOI-4507 b has a prograde orbit with $\lambda=-15_{-44}^{+50}$ deg. Although the reported error bars are relatively large, we show in Appendix C that the posteriors for the sky-projected obliquity are restricted between about -80 and 80 deg. Thus, the data strongly prefer a prograde orbit.
@@
$t_{0}$ Transit midpoint (BJD - 2,458,000) $\mathcal{U}(413.8472,413.8612)^{a}$ $413.8515\pm 0.0008$
@@
$\gamma_{\rm HARPS2}$ HARPS2 RV offset (m/s) $\mathcal{U}(25000,27000)$ $26095_{-10}^{+9}$
@@
VII Discussion
VII.1 TOI-4507 b in context
TOI-4507 b appears to be an interesting planet in several ways. Figure 4 shows the results for the projected obliquity $\lambda$ along with those of other systems. The obliquity data were drawn from TEPCat [^3] https://www.astro.keele.ac.uk/jkt/tepcat/ (J. Southworth, 2011) and the age data were drawn from the NASA Exoplanet Archive (R. L. Akeson et al., 2013; J. L. Christiansen et al., 2025), both as of May 2025. Evidently, TOI-4507 b has one of the longest orbital periods for which $\lambda$ has been measured, having the fourth-longest orbital period after HIP 41378 d (S. Grouffal et al., 2022), TIC 241249530 b (A. F. Gupta et al., 2024), and HD 80606 b (e.g., G. Hébrard et al., 2010). Here, we define Neptunes/sub-Saturns as the class of planets with $10<M_{p}/M_{\oplus}<90$, with the subclass of super-puffs having $\rho_{p}<0.3$ g/cm3.
TOI-4507 b joins a growing population of Neptune/sub-Saturn systems with measured stellar obliquities (e.g., V. Bourrier et al., 2023; J. I. Espinoza-Retamal et al., 2024; E. Knudstrup et al., 2024; L. B. Handley et al., 2025). This population includes some other super-puffs on prograde orbits, such as K2-261 b (E. Knudstrup et al., 2024), TOI-5398 b (G. Mantovan et al., 2024; B. T. Radzom et al., 2024), WASP-21 b (G. Chen et al., 2020), WASP-39 b (L. Mancini et al., 2018), WASP-69 b (N. Casasayas-Barris et al., 2017), WASP-117 b (M. Lendl et al., 2014), and WASP-193 b (S. W. Yee et al., 2025). Within this collection of prograde super-puffs, TOI-4507 b has the longest period.
@@ VII.1 TOI-4507 b in context
Additionally, with an age of $\sim 700$ Myr, TOI-4507 b is the youngest super-puff with an obliquity measurement. Further, it is one of the youngest Neptunes/sub-Saturns in general with an obliquity constraint, only behind AU Mic c (H. Yu et al., 2025), V1298 Tau c (A. D. Feinstein et al., 2021), and TOI-5398 b (G. Mantovan et al., 2024; B. T. Radzom et al., 2024).
@@ VII.2 Search for companions
If the RV trend is produced by a planetary companion, it should have an orbital period $P>5.6$ yr, which corresponds to the baseline of the current RV observations. The parameters of this possible companion could be constrained with additional RV and/or Gaia DR4 and DR5 astrometric measurements. Although TOI-4507 is relatively massive, it is nearby ($d\approx 177$ pc) and bright ($G\approx 10.6$), so it is a good candidate for Gaia astrometric detections (e.g., M. Perryman et al., 2014; C. Lammers & J. N. Winn, 2025). When removing the signal of TOI-4507 b and the RV trend from the data, we do not see clear additional signals in the RV data, which rules out the presence of planets producing similar or larger RV signals than TOI-4507 b.
@@
VII.3 Explaining the Inferred Low Density
@@
In summary, the primary challenge for the ring hypothesis is that long-lived opaque rings are unlikely to survive at the planet’s equilibrium temperature. If such rings were present, sustaining them in a high-obliquity configuration would require the planet to have a modified tidal quality factor of $Q_{p}/k_{2,p}\gtrsim 10^{5}$, a value that may be somewhat elevated but still plausible for a Neptune-like planet.
VIII Conclusions
We have reported the discovery and characterization of TOI-4507 b, a transiting cold Neptune on a prograde orbit. With a radius of $8.2\pm 0.1\,R_{\oplus}$, a mass $20\,M_{\oplus}$, and a bulk density <0.2<0.2 g/cm3, TOI-4507 b is among the longest-period ($P\sim 105$ d) super-puffs discovered to date. Further, TOI-4507 is one of the youngest Neptune/sub-Saturn systems with an obliquity constraint, and one of the longest-period systems in general for which the obliquity has been measured.
@@ VIII Conclusions
TOI-4507 thus joins a small but growing population of Neptune/sub-Saturn and super-puff systems with obliquity constraints, providing a valuable laboratory for understanding the formation, evolution, and diversity of planetary systems.
@@
We would like to thank Yubo Su for suggesting the rotational constraint on the sizes of the rings. We also thank Yugo Kawai for kindly pointing out that the source of the variability observed in the TESS light curve of TOI-4507 is a blended variable star in the background, and it is not related to the rotational period of TOI-4507. We would also like to thank the anonymous referee for their thoughtful review and suggestions that improved the quality of this work. We also thank Songhu Wang, Xian-Yu Wang, and the Red Worlds Lab group at the University of Amsterdam for their useful discussions and suggestions that strengthened this manuscript. J.I.E.-R. gratefully acknowledges support from the John and A-Lan Reynolds Faculty Research Fund, from the ANID BASAL project FB210003, and from the ANID Doctorado Nacional grant 2021-21212378. R.B. acknowledges support from FONDECYT Project 1241963 and from ANID – Millennium Science Initiative – ICN12_009. $T.$ acknowledges support by the BNSF program “VIHREN-2021" project No. KP-06-DV/5. Support for M.C. is provided by ANID’s Millennium Science Initiative through grants ICN12_12009 and AIM23-001, awarded to the Millennium Institute of Astrophysics (MAS); by ANID/FONDECYT Regular grant 1231637; and by ANID/Basal (CATA) grant FB21003. $C.$ acknowledges support from ANID through FONDECYT post-doctoral grant number 3230283. A.V.F. acknowledges the support of the Institute of Physics through the Bell Burnell Graduate Scholarship Fund. This paper was based on observations collected at the European Southern Observatory under ESO programmes 0104.C-0413(A), 106.21ER.001, 112.25W1.001, 114.27CS.001, and 115.286G.001. This paper also made use of data collected by the TESS mission and are publicly available from the Mikulski Archive for Space Telescopes (MAST) operated by the Space Telescope Science Institute (STScI). Funding for the TESS mission is provided by NASA’s Science Mission Directorate. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center for the production of the SPOC data products. This work makes use of observations from the ASTEP telescope. ASTEP benefited from the support of the French and Italian polar agencies IPEV and PNRA in the framework of the Concordia station program and from OCA, INSU, Idex UCAJEDI (ANR- 15-IDEX-01) and ESA through the Science Faculty of the European Space Research and Technology Centre (ESTEC). This research also received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 803193/BEBOP), from the Science and Technology Facilities Council (STFC; grant No. ST/S00193X/1, ST/W002582/1 and ST/Y001710/1), and from the ERC/UKRI Frontier Research Guarantee programme (CandY/ EP/Z000327/1). This work has been carried out within the framework of the National Centre of Competence in Research PlanetS supported by the Swiss National Science Foundation under grants 51NF40_182901 and 51NF40_205606. This paper is also based on observations obtained at the SOAR telescope, which is a joint project of the Ministério da Ciência, Tecnologia e Inovações (MCTI/LNA) do Brasil, the US National Science Foundation’s NOIRLab, the University of North Carolina at Chapel Hill (UNC), and Michigan State University (MSU). This research has made use of the Exoplanet Follow-up Observation Program (ExoFOP) and NASA Exoplanet Archive websites, which are operated by the California Institute of Technology under contract with NASA under the Exoplanet Exploration Program.